OpenStudy (anonymous):

how to find the first three derivatives of f(x)=2x^5 + x^(3/2) - (1/2x)?

OpenStudy (klimenkov):

You want to understand how or just to know the answer?

OpenStudy (anonymous):

Its just basic derivative....

OpenStudy (anonymous):

isn't it the 1st derivative for (-1/2x ) is (1)/(2x^2)? can you explain to me how did (-1/x^3) became the 2nd derivative for (-1/2x)?

OpenStudy (anonymous):

@klimenkov

OpenStudy (anonymous):

@Mitul

OpenStudy (anonymous):

For derivative, the rule is u go on decreasing the power of the variable and multiply that variable with that power each time u take derivative...

OpenStudy (klimenkov):

You are right. The first derivative for \(-\frac1{2x}\) is \(\frac{1}{2x^2}\). The second derivative is \(\left (\frac{1}{2x^2} \right) '\). Can you find it?

OpenStudy (anonymous):

see (-1/(2x)) means (-1/2)*x^(-1) first derivative : (-1/2)*-1*x^(-2) Second derivative : (-1/2)*-1*-2*(x^-3) and so on

OpenStudy (anonymous):

thanks..:)

OpenStudy (anonymous):

@Mitul how to find its third derivative?